OpenStudy (anonymous):

Solve the problem by stating the equation you would use to solve and then solving it. Midtown Antiques collects 5% sales tax on all sales. If total sales including tax are $1568.14, find the portion that is sales tax. Round to the nearest cent

OpenStudy (anonymous):

1568.14*0.05=78.40

OpenStudy (anonymous):

Divide by 105 Multiply by 100 Subtract this from initial sum ($1568.14)

OpenStudy (anonymous):

you want an equation right ? so lets say the initial cost is (i) and (t) is tax percentage (t)= .05 (i)= unknown so the total which is 1568.14 is (i) + (i * t)= 1568.14 we want to know (i * t) which is the portion that is sales tax but we don't know (i) so lets solve for (i) we can distribute (i) out (i)* ( 1 + (1 * t))= 1568.14 (i) * (1 + t ) = 1568.14 so (i) = 1568.14 / (1+t) (i) = 1568.14/ 1.05 now you know what (i) is so you can solve (i * t) now. Good Luck

OpenStudy (anonymous):

tks